A mixture of 5% disinfectant solution is to be made from 7% and 2% disinfectant solutions. How much of each solution should be used if 20 gallons of 5% solution are needed?
Answer by richwmiller(17219)

a+b=20,
0.07*a+0.02*b=0.05*20
a=20-b
0.07*(20-b)+0.02*b=1
1.4-0.07b+0.02*b=1
-0.05*b=-0.4
b=8
a=20-b
a=12 gal at 7%
b=8 gal at 2%
check
0.07*12+0.02*8=0.05*20
0.84+0.16=1
1=1
ok